After landing on an unfamiliar planet, a space explorer constructs a simple pendulum of length 45.0 cm . The explorer finds that the pendulum completes 106 full swing cycles in a time of 129 s .What is the magnitude of the gravitational acceleration on this planet?
Answer #1

Given that the pendulum completes 106 full swing cycles in 129s

Time Period of a simple pendulum is the time taken to complete 1 full oscillation.

Time Period = Total time/No of oscillations

T=\frac{129s}{106}

{\color{Red} T=1.217s}

=================

Use formula T= 271

\frac{T}{2\pi }=\sqrt{\frac{l}{g}}

(\frac{T}{2\pi })^{2}=\frac{l}{g}

\frac{T^{2}}{4\pi^{2} }=\frac{l}{g}

g=\frac{l*4\pi^{2} }{T^{2}}

g=\frac{0.45m*4\pi^{2} }{(1.217s)^{2}}

g=\frac{0.45*4\pi^{2} }{1.217^{2}}

ANSWER: {\color{Red} g=12m/s^{2}}

========================